# Corki: Enabling Real-time Embodied AI Robots via Algorithm-Architecture Co-Design

### v1

@@ 5.2. Baselines and Variations
We train RoboFlamingo using the Calvin dataset for accuracy comparison. The results are either higher or equivalent to the reported version. For latency and energy consumption comparisons, we establish a baseline using the traditional execution pipeline of embodied AI algorithms, where the inference latency, control latency, and communication latency are accumulated each frame.
@@
6. Evaluation
We evaluate Corki in this section. We first show Corki accelerator has a low hardware resource consumption (Sec. 6.1). We then evaluate both accuracy of Corki (Sec. 6.2) and corresponding latency and energy saving (Sec. 6.3).
6.1. Hardware Resource Consumption
The Corki accelerator is compact and does not require significant hardware resources, making it feasible for deployment on a real robot. It consumes only 13.6% of digital signal processors (DSP), 7.8% of flip-flops (FF), and 16.9% of look-up tables (LUT). The specialized on-chip buffer design is effective; the Corki accelerator utilizes only 6.6% of the total block random access memory (BRAM), with no data communication with off-chip DRAM during each control process.
6.2. Accuracy
@@
Fig. 11. Trajectory comparison between Corki and RoboFlamingo with two quantitative metrics.
@@
Fig. 12. Trajectory comparison of a randomly picked sequence from the test set. It is clearly shown that trajectories of Corki can follow the ground truth, while trajectories of Roboflamingo are off the target. We only show Corki-5 for simplicity.
@@
Among all fixed-step variations of Corki, Corki-5 achieves the highest accuracy and significantly outperforms the baseline. On seen tasks, it improves the average job length by 17.3% compared to the baseline, with a gain of 0.5 in job length. The trend observed among all Corki variations is that accuracy improves as the length of the actual trajectory taken increases. However, after reaching its peak accuracy, there is a gradual degradation in performance when the length of the actual trajectory taken continues to increase.
Corki-ADAP selects the length of the actual trajectory through identifying waypoints with significant movements. We observe that the results of Corki-ADAP fall between those of Corki-7 and Corki-5 in seen tasks, and it even outperforms Corki-5 in unseen tasks. This demonstrates that determining length during runtime is effective.
@@
The improvement brought by Corki is significant. Corki outperforms the baseline in almost all cases because trajectory naturally provides more robotic-friendly supervision during algorithm training. When the datasets of embodied AI algorithms are constructed, the collection of the ground truth was in the form of trajectory at first. In contrast, if discrete actions with 30 Hz frequency are used for supervision, the trajectory must first be decomposed into actions in a frame-basis and then used to train the model. Second, a smooth trajectory with high frequency control certainly improves success rate, which are demonstrated in robotic community (kleff2021high, ).
@@ 6.3. Performance Comparison
Fig. 13. Runtime latency and energy consumption comparison between Corki and baselines.
@@
We compare the latency and present the results in Fig. 13 on the left y-axis. Corki significantly reduces the frame latency of embodied AI robotic applications. Among the variations, Corki-9 achieves the best speedup of 3.6×3.6$\times$3.6 ×, as the inference frequency of the large language model is reduced by 8×8$\times$8 ×. As the length of the actual trajectory taken increases from 1 to 9, the speedup gradually increases from $\times1$ to 3.6×3.6$\times$3.6 ×. On the other hand, Corki-ADAP demonstrates a speedup of 3.0×3.0$\times$3.0 ×, providing an ideal trade-off between accuracy and efficiency.
@@
Corki also significantly saves energy consumption. Corki-1 has slightly higher energy consumption compared to the baseline, as it takes one step for every predicted trajectory, which is similar to the baseline. Besides Corki-1, all Corki variations have significantly lower energy consumption. Corki-9 has a 8.9×8.9$\times$8.9 × energy reduction. Low energy consumption is critical to robots, which are mostly battery-supported devices.
(a) Per-frame latency breakdown.
(b) Per-frame energy breakdown.
Fig. 14. Per-frame latency and energy comparison between Corki and RoboFlamingo.
@@
We finally show frame-by-frame analysis of latency and energy consumption for one single sequence. Fig. 14a shows the results of latency and Fig. 14b shows the results for energy consumption. Both latency and energy consumption of Corki are having a same trend, where the crest indicates the inference of LLM is happening at that time step, and trough means the robot is executing the trajectory predicted from the last time. Corki-5 has a periodical crest, as every 5 time steps, the inference will happen once. Corki-ADAP has a more flexible crest and trough, compared to Corki-5. This is due to the waypoints identification and flexible length of actual trajectory.
@@ 7. Related Work
Embodied Artificial Intelligence (EAI) differs from semantic AI by emphasizing agents, typically robots, that interact with the environment and execute long-horizon tasks. Recently, with the success of Large Language Models (LLMs) as planners, research in this domain has intensified, aiming to develop highly intelligent robots (duan2022survey, ; franklin1997autonomous, ; chrisley2003embodied, ; liu2024ok, ; vemprala2024chatgpt, ; huang2023voxposer, ). While most studies focus on enhancing functionalities, our research emphasizes real-time performance. Our approach is rooted in the robotic community, where trajectory serves as the fundamental unit of planning and control. This contrasts with the predominant vision-centric perspective, which treats images or frames as the basic units.
@@
With the growing interest in treating robots as a new computing platform, our community has increasingly focused on dedicated accelerators for robotic computing. These accelerators have been designed for localization (suleiman2019navion, ; gan2021eudoxus, ; liu2021archytas, ; liu2019eslam, ; eyvazpour2023hardware, ; sugiura2022universal, ; liu2022mobilesp, ; sugiura2021unified, ), motion planning (hsiao2023vapr, ; hao2023blitzcrank, ; huang2024moped, ; bakhshalipour2022racod, ; shah2023energy, ; lian2018dadu, ; murray2016microarchitecture, ; murray2019programmable, ), control (neuman2021robomorphic, ; neuman2023roboshape, ; yang2023dadu, ; lian2017dadu, ; sacks2018robox, ; shao2018towards, ; aude1991hardware, ; gac2012fpga, ), and more (mayoral2022robotcore, ; lienen2020reconros, ; hao2024orianna, ; lee2024spade, ; krishnan2022automatic, ; yu2020building, ). However, most accelerators focus on one or multiple modules within a traditional rule-based robotic computing system. Our work, in contrast, focuses on an end-to-end learning-based system, combining innovations in both algorithms and architecture, setting it apart from previous research.
@@ 8. Conclusion
Robots equipped with embodied AI algorithms often experience high latency due to the sequential execution pipeline and frequent LLM inference. In this paper, we propose Corki, a software-hardware co-design framework that significantly accelerates this process by transforming the algorithms to predict future trajectories, speeding up the control process, and pipelining communication with control. Results show that Corki achieves up to a 3.6×3.6$\times$3.6 × speedup. Corki also achieves a maximum 17.3% improvement in success rate.

### v5

@@ 5.2. Baselines and Variations
We train RoboFlamingo using the Calvin dataset for accuracy comparison. The results are either higher or equivalent to the reported version. For latency and energy consumption comparisons, we establish a baseline using the traditional execution pipeline of embodied AI algorithms, where the inference latency, control latency, and communication latency are accumulated in each frame.
@@
To demonstrate the effectiveness of the accelerator, we introduce a variant named Corki-SW, which employs trajectory prediction with five steps (as in Corki-5) while retaining the baseline CPU for control processing. It applies same approximation with Corki-5.
@@
Corki-SW (using Corki-5) 85.9% 68.4% 54.3% 42.2% 31.6% 2.824
6. Evaluation
We evaluate Corki in this section. We first show that the Corki accelerator has a low hardware resource consumption (Sec. 6.1). We then evaluate both the accuracy of Corki (Sec. 6.2) and corresponding latency and energy saving (Sec. 6.3).
6.1. Hardware Resource Consumption
The compact Corki accelerator requires minimal hardware resources, making it suitable for deployment on a real robot. It consumes only 13.6% of digital signal processors (DSP), 7.8% of flip-flops (FF), and 16.9% of look-up tables (LUT). The specialized on-chip buffer design is effective; the Corki accelerator utilizes only 6.6% of the total block random access memory (BRAM), with no data communication with off-chip DRAM during each control process.
6.2. Accuracy
@@
Figure 11. Trajectory comparison between Corki and RoboFlamingo with two quantitative metrics.
@@
Figure 12. Trajectory comparison of a randomly picked sequence from the test set. It is clearly shown that trajectories of Corki can follow the ground truth, while trajectories of Roboflamingo are off the target. We only show Corki-5 for simplicity.
@@
Among all fixed-step variations of Corki, Corki-5 achieves the highest accuracy and significantly outperforms the baseline. On seen tasks, the average job length is improved by 17.3% compared to the baseline, with a gain of 0.5 in job length. The trend observed among all Corki variations is that accuracy improves as the length of the actual trajectory taken increases. However, after reaching its peak accuracy, there is a gradual degradation in performance when the length of the actual trajectory taken continues to increase.
Corki-ADAP selects the length of the actual trajectory by identifying waypoints with significant movements. We observe that the results of Corki-ADAP fall between those of Corki-7 and Corki-5 in seen tasks, and it even outperforms Corki-5 in unseen tasks. This demonstrates that determining length during runtime is effective. Corki-SW achieves the same accuracy as Corki-5 because the only difference is whether the control process occurs on the accelerator, which does not affect accuracy.
@@
The improvement brought by Corki is significant. Corki outperforms the baseline in almost all cases because trajectory naturally provides a more robotic-friendly supervision during algorithm training. When the datasets of embodied AI algorithms are constructed, the collection of the ground truth was in the form of trajectory at first. In contrast, if discrete actions with 30 Hz frequency are used for supervision, the trajectory must first be decomposed into actions on a frame-basis and then used to train the model. Second, a smooth trajectory with high-frequency control certainly improves the success rate, which is demonstrated in the robotic community (Kleff et al., 2021).
@@ 6.3. Performance Comparison
Figure 13. Runtime latency and energy consumption comparison between Corki and baselines.
@@
We compare the latency and present the results in Fig. 13 on the left y-axis. Corki significantly reduces the frame latency of embodied AI robotic applications. Among the variations, Corki-9 achieves the best speedup of 9.1×9.1$\times$9.1 ×, as the inference frequency of the large language model is reduced by 8×8$\times$8 ×. As the length of the actual trajectory taken increases from 1 to 9, the speedup gradually increases from 1.2×1.2$\times$1.2 × to 9.1×9.1$\times$9.1 ×. On the other hand, Corki-ADAP demonstrates a speedup of 5.9×5.9$\times$5.9 ×, providing an ideal trade-off between accuracy and efficiency. Compared to Corki-5, Corki-SW has a 43.6% longer average latency, as the actual control still happens on CPUs. Specifically, Corki-5 has an average 26.9 Hz frequency, while Corki-SW only has 18.7 Hz.
@@
Corki also significantly saves energy consumption. Corki-1 has slightly higher energy consumption compared to the baseline, as it takes one step for every predicted trajectory, which is similar to the baseline. Besides Corki-1, all Corki variations have significantly lower energy consumption. Corki-9 has a 9.2×9.2$\times$9.2 × energy reduction. Low energy consumption is critical to robots, which are mostly battery-supported devices.
(a) Per-frame latency breakdown
(b) Per-frame energy breakdown
(c) Long tail problem and worst case latency analysis.
Figure 14. Per-frame latency, energy comparison and long-tail analysis.
@@
We finally show a frame-by-frame analysis of latency and energy consumption for one single sequence. Fig. 14a shows the results of latency, and Fig. 14b shows the results for energy consumption. Both latency and energy consumption of Corki have the same trend, where the crest indicates the inference of LLM happening at that time step, and the trough means the robot is executing the trajectory predicted from the last time. Corki-5 has a periodical crest, as every 5-time steps, the inference will happen once. Corki-ADAP has a more flexible crest and trough compared to Corki-5. This is due to the waypoints identification and flexible length of the actual trajectory.
We find that although our method achieves lower average frame latency, it does exhibit severer long tail problem, as different frames undergo different execution pipelines. We sort the frame latency and show the results in Fig. 14c, result suggests the relative latency variation of the baseline is 56.0% lower than Corki.
@@
We vary the baselines in Tbl. 3 and data representations in Tbl. 4 to better understand the performance. Our findings show that regardless of whether inference is performed on a state-of-the-art GPU like the H100 or an embedded GPU such as the Jetson Orin, Corki-ADAP consistently achieves high speedup (Jetson Orin shows longest inference latency, consuming over 0.9 second per frame, which makes it impossible to achieve real-time control). A similar conclusion holds when using different data representations.
Table 3. Performance under different GPU/CPU baselines.
GPUs V100 (Currently used) H100 Jetson Orin 32GB Xeon 8260
Normalized Inference Latency 1×1$\times$1 × 0.4×0.4$\times$0.4 × 10.0×10.0$\times$10.0 × 8.9×8.9$\times$8.9 ×
Speedup 5.9×5.9$\times$5.9 × 6.4×6.4$\times$6.4 × 5.3×5.3$\times$5.3 × 5.4×5.4$\times$5.4 ×
Table 4. Performance under different data representations.
GPUs 32-bit Float (Currently used) 16-bit Float 8-bit Int
Normalized Inference Latency 1×1$\times$1 × 0.8×0.8$\times$0.8 × 0.4×0.4$\times$0.4 ×
Speedup 5.9×5.9$\times$5.9 × 6.0×6.0$\times$6.0 × 6.4×6.4$\times$6.4 ×
6.4. Sensitivity Study
We further conduct experiments to analyze the impact of the approximation threshold on trajectory error and speedup within our control system. A higher threshold signifies a greater tendency to retain previously computed parameter values. Fig. 15a illustrates the relationship between the speedup and the approximation threshold. As the threshold increases, approximate computation becomes more prevalent, leading to an increase in the speedup. Fig. 15b depicts the relationship between the trajectory error and the threshold. Across the range of thresholds, the trajectory error remains minimal. This can be attributed to the reduction in computation latency afforded by approximate computation, allowing for higher control frequencies and consequently, more robust control. In our design, we opt for the threshold of $40\%$ to balance speedup and accuracy.
(a) Normalized speedup increases when the level of approximation increases.
(b) Trajectory error increases when the level of approximation increases.
Figure 15. Relationship between speedup and trajectory error with respect to the approximation threshold.
@@ 7. Related Work
Embodied Artificial Intelligence (EAI) differs from semantic AI by emphasizing agents, typically robots, that interact with the environment and execute long-horizon tasks. Recently, with the success of Large Language Models (LLMs) as planners, research in this domain has intensified, aiming to develop highly intelligent robots (Duan et al., 2022; Franklin, 1997; Chrisley, 2003; Liu et al., [n. d.]; Vemprala et al., 2024; Huang et al., 2023a). While most studies focus on enhancing functionalities, our research emphasizes real-time performance. Our approach is rooted in the robotic community, where trajectory serves as the fundamental unit of planning and control. This contrasts with the predominant vision-centric perspective, which treats images or frames as the basic units.
@@
With the growing interest in treating robots as a new computing platform, our community has increasingly focused on dedicated accelerators for robotic computing. These accelerators have been designed for localization (Suleiman et al., 2019; Gan et al., 2021; Liu et al., 2021, 2019; Eyvazpour et al., 2023; Sugiura and Matsutani, 2022; Liu et al., 2022b; Sugiura and Matsutani, 2021; Hao et al., 2022; Liu et al., 2022a), motion planning (Hsiao et al., 2023; Hao et al., 2023; Huang et al., 2024; Bakhshalipour et al., 2022; Shah et al., 2023; Lian et al., 2018; Murray et al., 2016, 2019), control (Neuman et al., 2021, 2023; Yang et al., 2023; Lian et al., 2017; Sacks et al., 2018; Shao et al., 2018; Aude and Aude, 1991; Gac et al., 2012), and more (Mayoral-Vilches et al., 2022; Lienen et al., 2020; Hao et al., 2024; Lee et al., 2024; Krishnan et al., 2022; Yu et al., 2020; Bakhshalipour and Gibbons, 2024). However, most accelerators focus on one or multiple modules within a traditional rule-based robotic computing system. Our work, in contrast, focuses on an end-to-end learning-based system, combining innovations in both algorithms and architecture, setting it apart from previous research.
8. Discussion
In this work, we focus on modifying the execution pipeline of embodied AI-powered robotic manipulation tasks, shifting from vision-centric frame-by-frame prediction to robot-centric trajectory prediction. Our results indicate that the proposed method performs well in a setting where a single robotic arm manipulates objects within a confined space, such as a desk.
Note that while the principle of predicting continuous trajectories instead of discrete actions can be extended to other robot types and tasks in the embodied AI domain, significant detailed design is required. At a minimum, we identify two key aspects to consider.
First, our method is limited to robotic arms, which typically have 9 DoF or fewer. Predicting the trajectory of a robot with a higher degree of freedom requires significantly greater effort. For instance, in the case of a humanoid robot, the trajectories of both the feet and arms must be predicted, while also considering their coordination. Simply applying our method to a humanoid robot may not work.
Second, our method can currently handle relatively long trajectories, given that sudden changes in the movement of a robotic arm are rare and that robotic arm’s motion tends to be slow. However, in tasks where the robot moves quickly with abrupt changes, the trajectory prediction must adapt accordingly.
Safety concerns.
This work focuses on slowly moving, space-constrained collaborative robotic arms, which were initially introduced as a safer alternative as opposed to industrial robots (Zanchettin et al., 2015). Compared to our baseline, the safety concern of our approach is predicting a longer future trajectory. We try to mitigate this risk by incorporating closed-loop features. On the other hand, the higher control frequency and lower trajectory errors demonstrate that Corki enables a much smoother, jitter-less control compared to the baseline, thereby reducing safety concerns during execution.
End-to-end system power.
The power and energy saving reported in this paper pertain solely to the computing system. If we include the energy consumed by the motors powering the robots, the overall energy savings would be lower. In our setting, the computing system inside the robot accounts for 40.6% of the total system power consumption (excluding server power).
@@ 9. Conclusion
Robots equipped with embodied AI algorithms often experience high latency due to the sequential execution pipeline and frequent LLM inference. In this paper, we propose Corki, a software-hardware co-design framework that significantly accelerates this process by transforming the algorithms to predict future trajectories, speeding up the control process, and pipelining communication with control. Results show that Corki achieves up to a 5.9×5.9$\times$5.9 × speedup. Corki also achieves a maximum 13.9% improvement in success rate.
Acknowledgements.
We thank the anonymous reviewers for their valuable feedback. This research was partially supported by the National Key Research and Development Program of China (Grant No. 2024YFB4505800), the Beijing Municipal Science and Technology Commission (Grant No. Z241100004224015) and the Longgang District Shenzhen’s “Ten Action Plan” for Supporting Innovation Projects (under Grant LGKCSDPT2024002), whose support is gratefully acknowledged. Feng Yan and Lin Ma are the first author’s mentors at Meituan.